OpenStudy (kodak.kiid):

Write 62 17/50 as a decimal number.

OpenStudy (anonymous):

is it 62 and 17/50ths or just 17/50ths?

Directrix (directrix):

62 17/50 17/50 = 34/100 = .34 62 + .34 = ? @Kodak.Kiid